Meeting Information

  • 1600 UT (1000 MDT)
  • AOC 317; IP: 146.88.16.8; phone: +1 575-835-7370

Agenda/Minutes

News, Travel, Visitors (McMullin)

  • Overall status: Last week of DNC (those that have data please review ASAP); projects are duking it out in the OST for time; double maintenance day Tuesday-Wednesday. An announcement will be added to operator logs asking for RFI monitoring feedback; this will be channeled through Dan/Gustaaf.
  • Key dates:
    • 01 Oct: Proposal deadline (A config; lowband for end of B config)
    • 04 Oct: Begin move to C configuration
    • 13-15 Oct: Phased Array testing campaign
  • Staff Travel:
    • Sowinski: 20 Sep - 15 Oct
    • Perley: 23 Sep - 19 Oct
    • McMullin: 01-05 Oct (vacation); 05-10 Oct (ALMA)
    • van Moorsel: 11-15 Oct
    • Wrobel: 02-09 Oct (ALMA); 24 Oct-27 Oct
    • DRAO: Dougherty: 11 Oct - 12 Oct (SKA; ABQ)
  • Visitors:
    • Brent Carlson: 18 October - 25 October; second trip post SKA meeting TBD
    • DRAO: Dougherty: 12 Oct - 17 Oct (NRAO-DSOC)
    • Adam Deller: 09 Oct - 16 Oct (phased array testing)
    • Joe Lazio: 12-15Oct (phased array testing)

Hardware status and updates (Langley, Perley, Rupen)

  • Array (Perley/Langley):
    • EA14 available tentatively on Thursday for science observations (has L, S, C, X (wide), Ku, K, Ka, Q); EA06 (L band - pushed to next week). Several RF switches have been replaced. Will be doing testing this week with rapid sampling to evaluate residual issues.
      • Antennas with 3-bit samplers: EA12, EA15, EA22, EA28
    • Links:
  • WIDAR (Rupen):
    • Notes: StB configuration queues (fell behind in setup; identified and fixed); BlB rebooted recently had their Chip IDs flipped resulting in the CBE discarding this data - this was the principal issue preventing RSRO Wednesday-Thursday last week; fix is going in today; post-3-bit observing resulted in 0s for most subbands; MR-generated script to run to reset for 8-bit operation; 3-bit testing status: still hunting down anomalously high noise characteristics-not yet understood.
  • Computing (Robnett):

Software status and updates (Butler)

Observing recommendations update (Ops)

Commissioning priorities for the week (McMullin, Chandler)

  • 0) Commissioning tests
  • 1) Science observing as possible
  • 2) Software development testing
  • Interrupt status for additional 3-bit testing

Commissioning Reports

Correlator: Momjian: Update: CBE Time Averaging Tests. https://bugs.nrao.edu/browse/ECSV-106

  • Update: 10s averaging works (some issues with the formed SDM; resolved in current release so an additional test to be run); application of flags to averaged (as well as non-averaged data) works on the basis of the time stamp for the given record (averaged or unaveraged); if the flag time range includes this time stamp, then it will be flagged if you apply it; if it ends before, then the data will not be flagged; some 'padding' of flagging times (a la quack) might be needed to properly handle the flagging.

Correlator: McMullin: MCAF update (compare test and production)

Planetary: Butler: Venus/TNO observations

Quick look data displays using TelCal files (Dhawan/McMullin)

In preparation

  • Polarization: Multi-band leakage/PA (Myers, Marvil): Not a priority but still needed; keep on list.
  • Correlator: Krauss/Hunter: 7 vs 8 node observing (follow-up; pending RSRO results)

Commissioning Test Planning

  • Array: RF switch tests with rapid monitoring
  • Correlator: CBE averaging with restricted nodes
  • Correlator: Quadrant 2 tests; CM update test.
  • Calibration: Switched power on/off 1dB attenuators.
  • Calibration: Switched power tracking over several airmasses.
  • Polarization: Antenna cross polarization (Perley; Need script.); LSCX on 3C147; Delayed until Rick's return
  • Spectral Line: Dhawan: Follow-up on frequency offset phase jumps characterization.
  • Spectral Line: Ott: Doppler tracking sampling follow-up
  • Spectral Line: Momjian: Zeeman test follow-up

RSRO/ECSO Reports

AK726: Kepley: DnC status; https://bugs.nrao.edu/browse/RSRO-6

AL746: Leroy: status; https://bugs.nrao.edu/browse/RSRO-7

10B-126: Takahashi: C band test run (TRSR0025) results. https://bugs.nrao.edu/browse/RSRO-31

  • Notes:
    • Current Residents
      • Volker Heesen (POC: Michael Rupen); area(s): polarimetry
      • Peter Hofner (POC: Juergen Ott); area(s): post-processing
      • Russ Taylor (POC: Rick Perley); area(s): polarimetry
      • Satoko Takahashi (POC: Juergen Ott); area(s): continuum, mosaicing; 10B-126 (AT380): A Millimeter Study of the Embedded Star-Cluster in the Orion Molecular Cloud.
        • Hauyu Liu (POC: Juergen Ott); area(s): continuum, mosaicing
      • Greg Hallinan (POC: Kumar Golap); area(s):; 10B-209 (AH1020): Broadband Periodic Dynamic Spectra of Ultracool Dwarf Pulsars
      • Adam Leroy (POC: TBD); area(s):; 10A-196 (AL746): Resolving the Starbursts in Nearby LIRGs and ULIRGs

Demo Science (Chandler/Condon): Project Codes: *TDEM*

Project Status Time Comments SB
Crab nebula Complete 10/10 L-band OSRO1/OSRO2 Complete
IRC+10216 Ready (HF) 21/33 Partial; revisit in DnC/C Partial Additional data taken!
M87 Ready (HF) 0/6 Ka Not done
M51 Ready (MR) 8/8   Complete
Cygnus A Ready 3x6 1x10 possible; 12 OSRO hours obtained Withdrawn
M82 Complete 2/2 L band Complete
Herc A Ready 3/5 CX bands Partial

Commissioning Group updates

AOB

Misc Actions

  • [McMullin/Frail]: %REDWrite up disposition of RR items.
  • [McMullin]: Confirm restriction of RSRO scan length at 90 s
  • [Owen]: Outline issue with 90s scan length restriction.
  • [McMullin/Durand]: Clean up outdated MainSaver issues.
  • [McMullin]: Review/recover metrics for Priority 0 items.
  • [Perley/McMullin]: SPW/IF/BB lexicon and mapper
  • List of known problems (can be used by OSRO commissioning team (van Moorsel)
      • Med Priority [Rupen]: X, Ku band subbands are inverted (IF 1 is labelled as IF4, IF2 as 3, IF3 as 2, IF 4 as 1). Noted that this had been tested but there were enough changes to warrant a re-do.
      • Med Priority [TBD]: Phase offsets within scans (phase jump at scan boundaries). R&L phase offset on some scans - probably L302 LO problem, needs tracking down. MR: Seen on EA07, EA25 (both RR and LL)
      • Med Priority [TBD]: Added. Phase change when frequency changes (needs to be tested).
      • Med Priority [TBD]: Anti-aliasing does not behave as advertised.
      • Low Priority [TBD]: First scan often bad.
      • Low Priority [TBD]: Phase (and amp) wobbles; long history.
-- JosephMcMullin - 22 Sep 2010
Topic revision: r3 - 2010-09-28, jmcmulli
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki? Send feedback